### v1

@@ V-B2 Time-Domain Analysis of Estimation Accuracy
Table IV presents the NMSE results segmented by time, distinguishing between the early and late portions of the signal (before and after $0.025$ s) at all the evaluation points. In the early segment, the proposed method improves the NMSE by approximately $16.6$ dB and $9.7$ dB relative to PINNs-tanh and PINNs-tanh-Res, respectively. Even compared with the strong baseline of PINNs-SIREN, an improvement of $1.0$ dB is achieved. In the late segment, the proposed method also maintains superior performance, improving the NMSE by approximately $7.4$ dB, $6.9$ dB, and $2.3$ dB compared with PINNs-tanh, PINNs-tanh-Res, and PINNs-SIREN, respectively. Overall, these results indicate that SIREN-based models provide strong accuracy through the early-reflection region, and that residual connections further improve the estimation of late-reverberation components.
VI Conclusion

### v2

@@ 5.2.2 Time-Domain Analysis of Estimation Accuracy
Table 4 presents the NMSE results segmented by time, distinguishing between the early and late portions of the signal (before and after $0.025$ s) at all the evaluation points. In the early segment, the proposed method improves the NMSE by approximately $15.0$ dB and $9.2$ dB relative to PINNs-tanh and PINNs-tanh-Res, respectively. Even compared with the strong baseline of PINNs-SIREN, an improvement of $0.4$ dB is achieved. In the late segment, the proposed method also maintains superior performance, improving the NMSE by approximately $6.8$ dB, $6.1$ dB, and $2.3$ dB compared with PINNs-tanh, PINNs-tanh-Res, and PINNs-SIREN, respectively. Overall, these results indicate that SIREN-based models provide strong accuracy through the early-reflection region, and that residual connections further improve the estimation of late-reverberation components.
6 Conclusion